# WorldLLM: Improving LLMs’ world modeling using curiosity-driven theory-making

### v1

@@ 3.2 WorldLLM with an oracle Experimenter
One can observe that, regardless of the oracle used or transition type, giving the Statistician hypotheses always improves its performance compared to when no hypotheses are given (i.e. the initial point of each curve). When comparing the oracles, one can observe that the data collection strategy of strongly influences WorldLLM’s abilities at producing efficient hypotheses. Indeed, using the same distribution as the one we perform evaluation on, O-Ideal achieves the best overall performance. On the opposite, O-Random obtains the poorest performance on complex transitions (i.e. growing an object) while being the best method on Standing transitions. This can easily be explained by the fact that moving actions are predominant in the action space. O-Hardest did not lead to strong performance, highlighting the need of diversity in collected transitions for the Scientist. Finally, O-Curriculum obtains results on par with O-Ideal. Results from Appendix D.2.1 indicate that O-Curriculum particularly helps the Scientist generating improved hypotheses.
@@ 3.3 WorldLLM with a curiosity-driven RL agent as Experimenter
As in the previous section, we analyze the evolution of the Statistician’s log-likelihoods over the test set. In order to ease the comparison of our three RL-based Experimenters to oracles, we introduce a coverage metric as the area under the evolution of log-likelihoods’ curve of each method. We normalize this area using the area between the Statistician’s log-likelihood without any hypotheses and the best performance achievable (i.e. 0). The complete formula is
[…]
The best performing method possible is expected to be very close to $1$ while an inept method’s result will be near 00. A method whose result is below 00 means that the hypotheses degrade the initial performance of the Statistician.
@@
Table 1: This table presents the normalized area under the training curve for each algorithm and transition type. The normalization is performed by dividing the computed area by the area formed by the initial log-likelihoods, i.e. obtained without using hypotheses.
From results shown in Table 1, one can observe that RL-ALPEXP is able to leverage its environment knowledge to reach performance significantly better than O-Random but remaining worse O-Ideal. When given no expert knowledge, RL-ALP achieves poor performance while RL-LogP obtains results close or even better than oracle baselines.
To better understand the effect of the RL-based Experimenters on the Statistician’s predictions, we studied the evolving distribution of collected transitions. We grouped Holding 1 and Holding 2 for simplicity and ignored Standing transitions which are predominant by nature. We show these results in Figure 4.
They show that RL-LogP manages to explore the environment and cover the whole transition space. The resulting collected distribution is near O-Hardest’s one (which is showed in dashed lines). This performance is notably explained by the syntax used by Playground-Text for observations and actions that makes complex transitions naturally harder to predict for an LLM (e.g. moving to a Big Herbivore involves only this entity whereas growing a Big Herbivore implies an action where two Plants are released on a baby Big Herbivore). This natural ordering appears in the rewards obtained by the RL policy and favors complex transitions. RL-ALPEXP, which does not leverage such a reward ordering, progressively converges to collecting more complex transitions. When looking at RL-ALP, most of the collected transitions are Standing ones. As we have seen in the previous section, the diversity of the transitions given to the Scientist is key in WorldLLM. Consequently, the hypotheses produced with RL-ALP focus solely on explaining Standing transitions. A more detailed analysis of why RL-ALP failed is available in Appendix D.4.1.
Figure 4: Proportion of collected transitions for RL-ALPEXP (left), RL-LogP (middle) and RL-ALP (right). The colored dashed lines correspond to the amount of transitions collected by O-Hardest.
This sections’ results show that WorldLLM successfully leads to improvement of the LLM’s forward modeling abilities when using curiosity-driven RL-based Experimenters. Moreover, we show that rewarding Experimenters for collecting transitions that help refine the hypotheses not only leads to efficient hypotheses but also leads to policies which solve our environment’s full technology tree.
@@ 3.4 Evaluating the produced world model
We now focus on evaluating the usefulness of the produced world model (i.e. the Statistician equipped with the last accepted hypotheses). While previous sections indicate that WorldLLM succeeds in increasing the likelihood of the observed next states, these results do not guarantee that the observed next states receive the highest likelihood compared to other possible next states. We thus examine our LLM-based world model’s capacity to correctly predict (i.e. generate) the next state on our test. In particular, for each state-action pair in $\mathcal{D}_{test}$, we used our Statistician’s LLM and constrained decoding to only generate the valid next states, given the current state, in Playground-Text. We then report in Table  2 the percentage of next states that belonged to the top-3 generated sequences.
@@ 3.5 Generating hypotheses with WorldLLM vs fine-tuning
We use the transitions collected by RL-LogP and fine-tune our Statistician’s LLM on them using a Causal Language Modeling objective (i.e. we maximize the likelihood of the next state’s tokens given the state-action pair). We name this new approach F-LogP. When evaluated on its top-3 constrained decoding accuracy over the test set as in Section 3.4, F-LogP obtains perfect results (i.e. accuracy of $1$) over all transition types. In the following sections, we study the generalization abilities of both methods and provide a deeper analysis of their impact on log-likelihoods.
3.5.1 Generalization to changes in the environment
We now assess whether using natural language hypotheses in the prompt produces better generalization abilities for the LLM-based world model than fine-tuning it on collected transitions. For this, we create a generalization environment in which the syntax of observations is changed (see Appendix 10). We use this new environment to report the top-3 constrained decoding accuracy on our test set as in previous sections.
@@ 3.5.1 Generalization to changes in the environment
Results from Table 3 show an important deterioration of the performance for all methods. F-LogP manages to maintain a non-zero accuracy on complex transition types while WorldLLM-based world models do not perform better than the Statistician without any hypotheses. These results hints that WorldLLM may struggle in finding abstract hypotheses that generalize. The next section provides additional insights on these results.
3.5.2 Effect on log-likelihoods
@@
When focusing on performance with the generalization environment, WorldLLM leads to very few changes compared to when no hypotheses are given. On its side, F-LogP produces very low log-likelihoods for all transitions and keeps the correct one slightly above the others. Overall these results show that both approaches generalize poorly.
One explanation for WorldLLM’s poor generalization performance lies in the hypotheses produced. We provided samples in Appendix D.3 which show that our Scientist notably used examples of transitions in the hypotheses (even though some hypotheses did mention abstract categories or even mentioned objects not in our environment). Such examples no longer align with the generalization environment’s syntax. Multiple explanations can be given for this. First, our current experiments rely on a Scientist’s LLM with limited capabilities. Second, while our objective in 1 minimizes the number of hypotheses, the current implementation of WorldLLM only focuses on maximizing the Statistician’s log-likelihood. We argue that favoring short set of hypotheses could help abstract theories emerge. We leave these investigations for future work.
(a) No changes to the environment.
(b) Generalization environment.
Figure 5: Log-likelihoods (averaged by transition type) produced by the Statistician when facing an instance of a Grow Plant transition. We show how using the generalization environment from Section 3.5.1 affects the log-likelihoods (b).
@@ 4 Related work
Building forward models
@@
Building forward models is a long-established research direction (Ha & Schmidhuber, 2018; Moerland et al., 2023). In addition to producing more interpretable agents, forward models were particularly used in RL to improve sample efficiency allowing the agent to train on imagined transitions or perform multi-step planning (Hafner et al., 2020; 2021; 2025; Schrittwieser et al., 2020). While these works learned the forward model at the same time as the RL policy, a line of work also studied how learning the forward model could be used as an intrinsic reward to foster the policy’s exploration (Schmidhuber, 1991b; Pathak et al., 2017; Burda et al., 2019a; Oudeyer et al., 2007). WorldLLM employs a similar curiosity-driven scheme where our RL-based Experimenters use reward signals from prior work to promote exploration and collect transitions helping to improve the forward model. However, our method differs in the way the model is improved: instead of using a gradient-based approach to fine-tune the LLM-based forward model, as done in (Xiang et al., 2023), WorldLLM searches for natural language theories that improve the forward model when given in the prompt. While recent work studied how a forward model could leverage external textual information such as a manual explaining the environment (Zhang et al., 2024; Dainese et al., 2023), our approach iteratively generates and refines its own explanations.
Inductive reasoning for world modeling
@@
The importance of theories in shaping one’s representations about the world was argued as central, in both children development and adults (Piaget, 1954; Gopnik, 1996). Building on this, the literature of theory-based RL has been advocating for augmenting artificial agents with inductive reasoning (Tsividis et al., 2021). In particular, (Tsividis et al., 2021; Tang et al., 2024a; Li et al., 2024) proposed approaches based on LLMs which generate programs representing the theories about an environment’s dynamics. The use of programs allowed the authors to execute and compare how the proposed theories match empirical evidence. Parallel to this, (Piriyakulkij et al., 2024; Wang et al., 2024) proposed to generate theories using natural language. While the former constrained the generated theories to programs which can be easily evaluated in their environment, the latter used the generated theories to condition a program generator to simulate the environment’s dynamics. WorldLLM also builds on the theory-based RL literature and closely relates to (Piriyakulkij et al., 2024). Notably, WorldLLM also leverages Bayesian Inference with an LLM as the proposal distribution to generate the theories. However, the major difference lies in how the produced theories are used: while (Piriyakulkij et al., 2024) constrain the format of theories such that they can be directly tested in the environment, our work focuses on giving these theories in another LLM’s prompt to improve its forward modeling capabilities. In other words, our work strongly differs in its perspective of producing theories that improve our LLM’s world modeling abilities instead of producing correct theories. Nonetheless, the Bayesian Inference framework we used in our Scientist is largely inspired by (Piriyakulkij et al., 2024). Both our work and theirs also consider an active learning paradigm to choose which experiments to perform for theories’ refinement. While their approach computes exact Information Gain of all possible experiments, ours relies on curiosity-driven RL agents and Information Gain approximates as rewards to overcome the intractability of exact Information Gain in most environments.
@@ 5 Discussion
We propose WorldLLM, a method to improve LLMs’ world modeling abilities by generating a set of natural language hypotheses that are given to an LLM when predicting the outcome of a state-action pair. Our method improves the LLM using human-interpretable theories with natural language while being less computationally costly than directly fine-tuning large-scale LLMs over collected data. WorldLLM is also based on an architecture that does not require any hand-crafted theory space. Our results showed that the adversarial framework between the Scientist and simple curiosity-driven RL Experimenters fuels an exploratory dynamic that covers the environment and allows the discovery of the underlying rules.
Nonetheless, our results show several limitations. First, the inner dynamics of Playground-Text are arguably simple. Moreover, the natural ordering created by the hardest transitions being longer and harder to predict is not expected in more complex environments. This could significantly affect RL-LogP’s results. Studying WorldLLM’s scaling to more complex environments (e.g. Mars (Tang et al., 2024b)) is therefore an important step. Secondly, our results were performed using a single LLM. This model showed its limitations on Playground-Text, especially in its capacity to generate high-level and abstract hypotheses when used in the Scientist. We believe that using a more efficient LLM in the Scientist in addition to taking into account the hypotheses length (e.g. adding a non-uniform prior in our proposal) should improve WorldLLM’s efficiency. Then, our paper studied three simple curiosity-driven RL Experimenters. Studying more advanced methods such as autotelic approaches (Colas et al., 2022) could improve the quality of the collected transitions, which is key in our framework. Finally, our online Metropolis approach for the Scientist suffers from several limitations. Notably, it only considers the transitions collected at time $t$, creating potential forgetting when the Experimenter is non-stationary. This choice is motivated by the fact that 1) our LLM used as the proposal has a fixed and limited context window and thus cannot consider all transitions collected up to time $t$, and 2) computing the likelihood on a large set of transitions is costly. This could be improved by either keeping track of past data (e.g. using a buffer containing transitions from different timesteps) or considering more advanced Bayesian inference methods more suited for online data such as particle filters. The latter would also overcome another current limitation of our approach: a single particle is used and may lead our Scientist to get stuck in local optima (as hinted by Appendix D).
@@ Appendix A The Playground-Text Environment
We use the Playground-Text environment from (Gaven et al., 2024), itself being a textual adaptation of the Playground environment introduced by (Colas et al., 2020). This textual environment returns, at each time step, a description of its state and possible actions. In our experiments, we used a modified version of Playground-Text where the current state still describes the whole environment’s state but the next state only decribes how the performed action changed the environment (see Figure 6).
Playground-Text features 4 different classes of objects with different varieties: Water, Plants (carrot, potato, beet, berry and pea), Small Herbivores (pig, cow and sheep) and Big Herbivores (elephant, giraffe, rhinoceros). As depicted in Figure 2, all objects except Water possess two states: young (seed for plants and baby for animals) and grown. These objects transition from their young state to their grown one by consuming other objects.

### v2

@@ 3.2 WorldLLM with an oracle Experimenter
One can observe that, regardless of the oracle used or transition type, giving the Statistician hypotheses always improves its performance compared to when no hypotheses are given (i.e., the initial point of each curve). When comparing the oracles, one can observe that the data collection strategy of O-Ideal strongly influences WorldLLM’s abilities at producing efficient hypotheses. Indeed, using the same distribution as the one we perform evaluation on, O-Ideal achieves the best overall performance. On the opposite, O-Random obtains the poorest performance on complex transitions (i.e., growing an object) while being the best method on Standing transitions. This can easily be explained by the fact that moving actions are predominant in the action space. O-Hardest did not lead to strong performance, highlighting the need of diversity in collected transitions for the Scientist. Finally, O-Curriculum obtains results that are on par with those of O-Ideal. Results from Appendix D.2.1 indicate that O-Curriculum particularly helps the Scientist generate improved hypotheses.
@@ 3.3 WorldLLM with a curiosity-driven RL agent as Experimenter
As in the previous section, we analyze the evolution of the Statistician’s log-likelihoods over the test set. In order to ease the comparison of our three RL-based Experimenters to oracles, we introduce a coverage metric as the area under the evolution of log-likelihoods’ curve of each method. We normalize this area using the area between the Statistician’s log-likelihood without any hypotheses and the best performance achievable (i.e., 0). The complete formula is
$\text{area}=1-\mathop{\text{$\int_{\text{$\scriptstyle\kern-0.73497pt0$}}^{\text{$\scriptstyle T$}}$}}\nolimits\frac{\log\mathbb{P}^{S_{t}}(\mathcal{D}_{test}|H_{t})}{\log\mathbb{P}^{S_{t}}(\mathcal{D}_{test}|\{\emptyset\})T}dt.$ (3)
The best performing method possible is expected to be very close to $1$ while an inept method’s result will be near 0. A method whose result is below 0 means that the hypotheses degrade the initial performance of the Statistician.
@@
Table 1: This table presents the normalized area under the training curve for each algorithm and transition type. The normalization is performed by dividing the computed area by the area formed by the initial log-likelihoods, i.e., obtained without using hypotheses.
From the results shown in Table 1, one can observe that RL-ALPEXP is able to leverage its environment knowledge to reach performance significantly better than O-Random but remaining worse than O-Ideal. When given no expert knowledge, RL-ALP achieves poor performance while RL-LogP obtains results close to or even better than oracle baselines.
To better understand the effect of the RL-based Experimenters on the Statistician’s predictions, we studied the evolving distribution of collected transitions. We grouped Holding 1 and Holding 2 for simplicity and ignored Standing transitions, which are predominant by nature. We show these results in Figure 4.
They show that RL-LogP manages to explore the environment and cover the whole transition space. The resulting collected distribution is near O-Hardest’s one (which is shown in dashed lines). This performance is notably explained by the syntax used by Playground-Text for observations and actions that make complex transitions naturally harder to predict for an LLM (e.g., moving to a Big Herbivore involves only this entity whereas growing a Big Herbivore implies an action where two Plants are released on a baby Big Herbivore). This natural ordering appears in the rewards obtained by the RL policy and favors complex transitions. RL-ALPEXP, which does not leverage such a reward ordering, progressively converges to collecting more complex transitions. When looking at RL-ALP, most of the collected transitions are Standing ones. As we have seen in the previous section, the diversity of the transitions given to the Scientist is key in WorldLLM. Consequently, the hypotheses produced with RL-ALP focus solely on explaining Standing transitions. A more detailed analysis of why RL-ALP failed is available in Appendix D.4.1.
Figure 4: Proportion of collected transitions for RL-ALPEXP (left), RL-LogP (middle), and RL-ALP (right). The colored dashed lines correspond to the amount of transitions collected by O-Hardest.
This section’s results show that WorldLLM successfully improves the LLM’s forward modeling abilities when using curiosity-driven RL-based Experimenters. Moreover, we show that rewarding Experimenters for collecting transitions that help refine the hypotheses not only leads to efficient hypotheses but also leads to policies that solve our environment’s full technology tree.
@@ 3.4 Evaluating the produced world model
We now focus on evaluating the usefulness of the produced world model (i.e., the Statistician equipped with the last accepted hypotheses). While previous sections indicate that WorldLLM succeeds in increasing the likelihood of the observed next states, these results do not guarantee that the observed next states receive the highest likelihood compared to other possible next states. We thus examine our LLM-based world model’s capacity to correctly predict (i.e., generate) the next state on our test. In particular, for each state-action pair in $\mathcal{D}_{test}$, we used our Statistician’s LLM and constrained decoding to only generate the valid next states, given the current state, in Playground-Text. We then report in Table  2 the percentage of next states that belonged to the top-3 generated sequences.
@@ 3.5 Generating hypotheses with WorldLLM vs fine-tuning
We use the transitions collected by RL-LogP and fine-tune our Statistician’s LLM on them using a causal language modeling objective (i.e., we maximize the likelihood of the next state’s tokens given the state-action pair). We name this new approach F-LogP. When evaluated on its top-3 constrained decoding accuracy over the test set as in Section 3.4, F-LogP obtains perfect results (i.e., accuracy of $1$) over all transition types. In the following sections, we study the generalization abilities of both methods and provide a deeper analysis of their impact on log-likelihoods.
Generalization to changes in the environment
We now assess whether using natural language hypotheses in the prompt produces better generalization abilities for the LLM-based world model than fine-tuning it on collected transitions. For this, we create a generalization environment in which the syntax of observations is changed (see Figure 10). We use this new environment to report the top-3 constrained decoding accuracy on our test set, as in previous sections.
@@
Results from Table 3 show an important deterioration of the performance for all methods. F-LogP manages to maintain a non-zero accuracy on complex transition types while WorldLLM-based world models do not perform better than the Statistician without any hypotheses. These results hint that WorldLLM may struggle in finding abstract hypotheses that generalize. The next section provides additional insights on these results.
Effect on log-likelihoods
@@
When focusing on performance with the generalization environment, WorldLLM leads to very few changes compared to when no hypotheses are given. On its side, F-LogP produces very low log-likelihoods for all transitions and keeps the correct one slightly above the others. Overall, these results show that both approaches generalize poorly.
One explanation for WorldLLM’s poor generalization performance lies in the hypotheses produced. We provided samples in Appendix D.3 which show that our Scientist notably used examples of transitions in the hypotheses (even though some hypotheses did mention abstract categories or even mentioned objects not in our environment). Such examples no longer align with the generalization environment’s syntax. Multiple explanations can be given for this. First, our current experiments rely on a Scientist’s LLM with limited capabilities. Second, while our objective in 1 minimizes the number of hypotheses, the current implementation of WorldLLM only focuses on maximizing the Statistician’s log-likelihood. We argue that favoring short hypotheses could help abstract theories emerge. We leave these investigations for future work.
Figure 5: Log-likelihoods (averaged by transition type) produced by the Statistician when facing an instance of a Grow Plant transition. We show how using the generalization environment from Section 3.5 affects the log-likelihoods (b).
@@ 4 Related work
4.1 Building forward models
@@
Building forward models is a long-established research direction (ha_world_2018; moerland_model-based_2022). In addition to producing more interpretable agents, forward models were particularly used in RL to improve sample efficiency, allowing the agent to train on imagined transitions or perform multi-step planning (hafner_dream_2020; hafner_mastering_2022; hafner_mastering_2024; schrittwieser_mastering_2020). Parallel to this, learning a forward model has been used as an intrinsic reward to foster RL policies’ exploration (schmidhuber_possibility_1991; pathak_curiosity-driven_2017; burda_large-scale_2018; oudeyer_what_2007). WorldLLM employs a similar curiosity-driven scheme where our RL-based Experimenters use reward signals from prior work to promote exploration and collect transitions that help improve the forward model. However, our method differs in the way the model is improved: instead of using a gradient-based approach to fine-tune the LLM-based forward model, as done in (xiang_language_2023), WorldLLM searches for natural language theories that improve the forward model when given in the prompt. While recent work studied how a forward model could leverage external textual information such as a manual explaining the environment (zhang_language-guided_2024; dainese_reader_2023), our approach iteratively generates and refines its own explanations.
4.2 Inductive reasoning for world modeling
@@
The importance of theories in shaping one’s representations about the world was argued as central, in both children development and adults (piaget_construction_1954; gopnik_scientist_1996) Building on this, the literature of theory-based RL has been advocating for augmenting artificial agents with inductive reasoning (tsividis_human-level_2021). In particular, (tsividis_human-level_2021; tang_worldcoder_2024; li_automated_2024) proposed approaches based on LLMs that generate programs representing the theories about an environment’s dynamics. The use of programs allowed the authors to execute and compare how the proposed theories match empirical evidence. Parallel to this, (piriyakulkij_doing_2024; wang_hypothesis_2023) proposed to generate theories using natural language. While the former constrained the generated theories to programs that can be easily evaluated in their environment, the latter used the generated theories to condition a program generator to simulate the environment’s dynamics. WorldLLM also builds on the theory-based RL literature and closely relates to (piriyakulkij_doing_2024). Notably, WorldLLM also leverages Bayesian inference with an LLM as the proposal distribution to generate the theories. However, the major difference lies in how the produced theories are used: while (piriyakulkij_doing_2024) constrain the format of theories such that they can be directly tested in the environment, our work focuses on giving these theories in another LLM’s prompt to improve its forward modeling capabilities. In other words, our work strongly differs in its perspective of producing theories that improve our LLM’s world modeling abilities instead of producing correct theories. Nonetheless, the Bayesian inference framework we used in our Scientist is largely inspired by (piriyakulkij_doing_2024). Both our work and theirs also consider an active learning paradigm to choose which experiments to perform for theories’ refinement. While their approach computes exact Information Gain of all possible experiments, ours relies on curiosity-driven RL agents and Information Gain approximates as rewards to overcome the intractability of exact Information Gain in most environments.
@@ 5 Discussion
In this paper, we introduced WorldLLM, a framework for enhancing world modeling abilities in LLMs. WorldLLM generates natural language hypotheses—interpretable, theory-like statements—which condition the LLM when predicting state-action outcomes. This approach enables world modeling without expensive fine-tuning of the LLM itself and avoids the need for a handcrafted theory space. Our experiments demonstrated that the adversarial dynamic between a Scientist (hypothesis generator) and simple curiosity-driven Experimenters (data collectors) supports broad exploration and the discovery of environment dynamics.
Despite promising results, our initial implementation of WorldLLM presents several limitations. First, the Playground-Text environment is structurally simple, and the natural ordering created by the most challenging transitions being longer and harder to predict is not expected in more complex environments—potentially reducing the effectiveness of RL-LogP. Evaluating WorldLLM in more complex environments is therefore a critical direction for future work (ying_assessing_2025).
Second, our current Experimenters use lightweight neural networks and symbolic inputs. A more compelling alternative would be to integrate LLM-based Experimenters, possibly unifying all WorldLLM components with a single GLAM-equipped LLM that learns both to control and to improve its forward model via online RL. Furthermore, our experiments only used one LLM model, which showed limitations in the Scientist role—particularly in generating abstract, high-level hypotheses.
In addition to scaling the LLM, we argue that the emergence of such abstract hypotheses may result from other modifications to our implementation. A key advantage of Bayesian inference in this context is the ability to incorporate priors, which constrain the theory space. While a length-based prior might encourage abstract hypotheses, we believe richer priors—such as “error maps” that represent structured patterns of predictive failure, as described by schulz_chapter_2012—may be necessary to better replicate human-like theory refinement.
Additionally, our current Metropolis-based approach for updating hypotheses only considers the most recent transitions, which can lead to forgetting, especially with non-stationary Experimenters. This was primarily motivated by practical constraints: the LLM’s limited context window and the cost of evaluating likelihoods on large transition sets. These issues could be mitigated by storing past transitions in a replay buffer or employing more sophisticated Bayesian inference methods (e.g., particle filters), which would also allow tracking multiple particles to avoid local optima—a current limitation evidenced in Appendix D.
Finally, our exploration strategy relied solely on simple curiosity-driven signals focused on poorly predicted transitions. Other intrinsic motivation mechanisms have been studied in both humans and artificial agents. A particularly important one is autotelic learning, where learners generate, select, and attempt to master self-defined goals (steels_autotelic_2004; colas_autotelic_2022). Studying the use of such more advanced curiosity signals could significantly improve the quality of collected transitions.
@@ Appendix A The Playground-Text Environment
We use the Playground-Text environment from (gaven2024sacglamimprovingonlinerl), itself being a textual adaptation of the Playground environment introduced by (colas_language_2020). This textual environment returns, at each time step, a description of its state and possible actions. In our experiments, we used a modified version of Playground-Text where the current state still describes the whole environment’s state, but the next state only describes how the performed action changed the environment (see Figure 6).
Playground-Text features 4 different classes of objects with different varieties: Water, Plants (carrot, potato, beet, berry, and pea), Small Herbivores (pig, cow, and sheep), and Big Herbivores (elephant, giraffe, rhinoceros). As depicted in Figure 2, all objects except Water possess two states: young (seed for plants and baby for animals) and grown. These objects transition from their young state to their grown one by consuming other objects.